OpenStudy (anonymous):

a team has won 60 percent of the 20 games it has playes so far this season. If the team plays a total of 50 games all season and wins 80 percent of the remaining games, then the percent of games it won for the entire season will be 1. 66% 2. 69% 3. 70% 4. 72% 5. 74%

OpenStudy (a_clan):

So far team has won =(60/100)*20 = 12 games Total games played in season =50 80 percent of the remaining games won = 80 percent of (50-20) = (80/100)*30 = 24 games. Total game won this season = 36 So, percent games won in the entire season = (36/50)*100=72% Answer (4)